# SAIP-Net: Enhancing Remote Sensing Image Segmentation via Spectral Adaptive Information Propagation

### v1

@@
Abstract.
Semantic segmentation of remote sensing imagery demands precise spatial boundaries and robust intra-class consistency, challenging conventional hierarchical models. To address limitations arising from spatial domain feature fusion and insufficient receptive fields, this paper introduces SAIP-Net, a novel frequency-aware segmentation framework that leverages Spectral Adaptive Information Propagation. SAIP-Net employs adaptive frequency filtering and multi-scale receptive field enhancement to effectively suppress intra-class feature inconsistencies and sharpen boundary lines. Comprehensive experiments demonstrate significant performance improvements over state-of-the-art methods, highlighting the effectiveness of spectral-adaptive strategies combined with expanded receptive fields for remote sensing image segmentation.
@@ Abstract.
Figure 1. Overview of the challenges and motivations behind SAIP-Net. Remote sensing images often exhibit large intra-class variance, small inter-class differences, and irregular class layouts. To address these challenges, SAIP-Net uses the combination of four modules to improve intra-class consistency and enhance boundary accuracy, significantly improving the segmentation of complex structures in remote sensing images.
1. Introduction
@@
Figure 2. Overview of the proposed SAIP-Net architecture. The network combines: a Learnable High-Pass Filter Stem to enhance edge details, a Transformer Encoder that extracts global context via multi-stage learnable pooling, and a Frequency-Aware Decoder that fuses high- and low-level features using SAFF modules, while integrated Composite Dilated Convolution Layers expand the receptive field. These modules ultimately lead to result in improved intra-class consistency and enhanced boundary accuracy, thereby significantly improving the segmentation of complex structures in remote sensing images.
@@ 2. Related Work
Semantic segmentation assigns semantic labels to each pixel and is widely applied in both natural and remote sensing images. While the two domains share foundational techniques, their data characteristics and task constraints differ significantly. Natural image segmentation focuses on perspective-view scenes with diverse objects and consistent contexts (Long et al., 2015), while remote sensing imagery features large-scale, top-down views with high intra-class variability, inter-class similarity, which requires domain-specific solutions (Zhu et al., 2017).
2.1. Semantic Segmentation in Natural Images
@@
Early methods, such as thresholding and region growing, were limited by image quality and computational constraints. The advent of deep learning, particularly Convolutional Neural Networks (CNNs) (LeCun et al., 1998) and Fully Convolutional Networks (FCNs) (Long et al., 2015), enabled end-to-end pixel-level prediction, forming the basis for encoder-decoder architectures like U-Net (Ronneberger et al., 2015) and SegNet (Badrinarayanan et al., 2017). These models extract high-level semantic features while preserving spatial resolution.
@@
Subsequent research addressed challenges such as class imbalance, small object detection, and complex backgrounds. Dilated convolutions (Li et al., 2018) expanded receptive fields, while PSPNet (Zhou et al., 2019) introduced pyramid pooling to aggregate multi-scale contextual information. Attention mechanisms, such as Coordinate Attention (Hou et al., 2021), further refine segmentation, especially for small or complex objects. Additionally, unsupervised and semi-supervised approaches (Ahn et al., 2019; Lin et al., 2024) have been explored to reduce reliance on large annotated datasets.
2.2. Semantic Segmentation in Remote Sensing Images
@@
Remote sensing segmentation faces unique challenges, including large-scale scenes, spectral diversity, and subtle intra-class variation. While FCN-based methods (Chen et al., 2018; Sun et al., 2019; Li et al., 2021; Xue et al., 2022; Ma et al., 2022; Hou et al., 2022) remain prevalent, additional modules have been developed to enhance spatial and contextual feature extraction. Techniques like channel attention (Hu et al., 2018) and region-aware methods, such as S-RA-FCN (Mou et al., 2020) and HMANet (Niu et al., 2021), address feature redundancy and spatial dependencies. Edge-aware strategies, like Edge-FCN (He et al., 2020), improve boundary precision by incorporating edge detection.
@@
Recently, Transformer-based models (Wang et al., 2022a; Xu et al., 2023; Xie et al., 2021; Wang et al., 2022b; Chen et al., 2021; Sun et al., 2022) have gained attention for their ability to model long-range dependencies via self-attention. For instance, RSSFormer (Xu et al., 2023) introduces adaptive feature fusion to reduce background noise and enhance object saliency. However, Transformers still face challenges in capturing fine local details and accurately delineating object boundaries.
2.3. Frequency-domain Methods for Semantic Segmentation
@@
Frequency-domain feature processing has emerged as a promising direction in semantic segmentation. Shan et al. decomposed images via the Fourier transform into high and low frequency components (Shan et al., 2021), which capture edge and body information respectively, ensuring object consistency and edge supervision through deep fusion. Zhang et al. proposed FsaNet (Zhang et al., 2023), leveraging a frequency self-attention mechanism on low-frequency components to drastically reduce computational costs. Chen et al. introduced Frequency-Adaptive Dilated Convolution (FADC) (Chen et al., 2024b), dynamically adjusting dilation rates by local frequency patterns and enhancing feature bandwidth/receptive fields via adaptive kernels and frequency selection. Their FreqFusion (Chen et al., 2024a) further improves boundary sharpness and feature coherence for dense prediction tasks. Ma et al. developed AFANet (Ma et al., 2025), optimizing semantic structures with a CLIP-guided spatial adapter to boost weakly-supervised few-shot segmentation. These advances underscore the transformative potential of frequency-aware strategies in semantic segmentation.
@@
3. Methods
@@
3.1. Transformer Encoder
Our transformer encoder is designed to balance global context modeling and computational efficiency via pooling attention and hierarchical feature extraction. The encoder consists of $4$ stages. Following (Li et al., 2022), each stage of our encoder includes multiple Transformer blocks, and spatial downsampling between stages is achieved through learnable pooling operations.
@@ 3.1. Transformer Encoder
(1) $Q=\mathcal{P}_{Q}(X_{in}W_{Q}),\quad K=\mathcal{P}_{K}(X_{in}W_{K}),\quad V=\mathcal{P}_{V}(X_{in}W_{V}),$
@@
(2) $X_{l}=\mathrm{Attn}(Q,K,V)+Q=\mathrm{Softmax}\left(\frac{QK^{\top}+E^{(\mathrm{rel})}}{\sqrt{D}}\right)V+Q$
where $E^{(\mathrm{rel})}{ij}=Q_{i}\cdot R{p(i),p(j)}$. Here, $Q_{i}$ represents the query vector at token position $i$, and $R_{p(i),p(j)}$ denotes the learnable relative positional embedding between tokens $i$ and $j$, where $p(i)$ and $p(j)$ map token indices to their spatial locations. The inner product $Q_{i}\cdot R_{p(i),p(j)}$ injects relative position bias into the attention logits. To reduce the number of learnable parameters, we follow (Li et al., 2022) and decompose $R_{p(i),p(j)}$ along spatial dimensions as:
(3) $R_{p(i),p(j)}=R^{\mathrm{h}}_{h(i),h(j)}+R^{\mathrm{w}}_{w(i),w(j)},$
@@
3.2. Spectral Adaptive Feature Fusion Module
Figure 3. (a). Illustration of Spectral Adaptive Feature Fusion (SAFF) module. The module integrates high-level and low-level features using content guided low-pass and high-pass filters alongside spatial offset estimation. The structure of $\mathcal{F}^{\text{HP}}$ is enlarged at the bottom while $\mathcal{F}^{\text{LP}}$ shares a similar architecture. (b). Our module enhances intra-class consistency and boundary accuracy at the feature level, leading to better segmentation results.
@@ 3.2. Spectral Adaptive Feature Fusion Module
Inspired by (Chen et al., 2024a; He et al., 2012; Ma et al., 2019), incorporating a guidance signal (e.g., low-level features) provides valuable local structural priors. These priors help direct the filtering process such that the spatially varying kernels can adapt based on local content, ensuring that essential details like edges and textures are accurately preserved while undesired noise is suppressed. Additionally, guided filters that vary at each position allow fine-grained feature reassembly and precise spatial alignment, ensuring that the fused features accurately reflect both the global context and the local details. Such flexibility directly contributes to improved overall segmentation quality and enhanced boundary delineation.
@@
(4) $\tilde{\mathbf{Y}}^{l+1}=\mathcal{F}^{\text{LP}}(X^{l},\mathbf{Y}^{l+1}),\quad\tilde{\mathbf{X}}^{l}=\mathcal{F}^{\text{HP}}(X^{l},\mathbf{X}^{l})+\mathbf{X}^{l},\quad\mathbf{Z}^{l}=\tilde{\mathbf{Y}}^{l+1}+\tilde{\mathbf{X}}^{l}.$
@@
(5) $\hat{\mathbf{X}}^{l}=\mathcal{F}^{\text{HP}}(Z^{l},\mathbf{X}^{l})+\mathbf{X}^{l},\quad\hat{\mathbf{Y}}^{l+1}=\mathcal{F}^{\text{LP}}(Z^{l},\mathbf{Y}^{l+1}).$
@@
(6) $u(i,j),v(i,j)=\mathcal{OG}(Z^{l})_{ij},\quad\mathbf{Y}^{l}_{i,j}=\hat{\mathbf{Y}}^{l+1}_{i+u(i,j),j+v(i,j)}+\hat{\mathbf{X}}^{l}_{i,j}.$
@@
The key components of this module are two adaptive filter: Content Guided Low-Pass/High-Pass Filter $\mathcal{F}^{\text{LP}}$/$\mathcal{F}^{\text{HP}}$ and an offset generator $\mathcal{OG}$. We elaborate on each component as follows:
@@
Inspired by (He et al., 2012; Ma et al., 2019), we take $\mathbf{Z}$ as the guidance feature to predict the spatially variant filter weights filtering feature map $\mathbf{X}$. The filter consists of several steps:
  1. (1)
@@
    $\mathbf{\bar{V}}=\text{Conv}_{3\times 3}\!\bigl{(}\mathbf{Z}\bigr{)},\quad\mathbf{\bar{W}}^{p,q}_{i,j}=\begin{cases}\frac{\exp(\mathbf{\bar{V}}^{p,q}_{i,j})}{\sum_{p,q\in\Omega}\exp(\mathbf{\bar{V}}^{p,q}_{i,j})},&\text{(low-pass)}\\ \mathbf{E}\;-\;\frac{\exp(\mathbf{\hat{V}}_{i,j}^{p,q})}{\sum_{p,q\in\Omega}\exp(\mathbf{\hat{V}}_{i,j}^{p,q})},&\text{(high-pass)}\\ \end{cases}$
@@
    (8) $\mathbf{Y}^{g}_{i,j}=\sum_{p,q\in\Omega}\mathbf{\bar{W}}^{g,p,q}_{i,j}\,\mathbf{X}_{i+p,\,j+q},$
@@
  3. (3)
    Upsampling of High-Level Features. If the input is high-level features, the four groups $\mathbf{Y^{g}}$ are rearranged using Pixel Shuffle(Shi et al., 2016) to obtain a smooth 2×2$\times$2 × upsampled feature map $\mathbf{Y}\in\mathbb{R}^{C\times 2H\times 2W}$.
@@
    (9) $\mathbf{S}^{p,q}_{i,j}=\frac{\sum_{c}\mathbf{Z}_{c,i,j}\,\mathbf{Z}_{c,i+p,j+q}}{\sqrt{\sum_{c}(\mathbf{Z}_{c,i,j})^{2}}\;\sqrt{\sum_{c}(\mathbf{Z}_{c,i+p,j+q})^{2}}},$
  2. (2)
@@
    $\mathbf{D}=\text{Conv}_{3\times 3}([\mathbf{Z},\mathbf{S}]),\quad\mathbf{A}=\sigma\!\bigl{(}\text{Conv}_{3\times 3}([\mathbf{Z},\mathbf{S}])\bigr{)},$
    Where []delimited-[][...][ … ] denotes the concat operation and $\sigma$ denotes the sigmoid function. Finally, the offset is given by $\mathbf{O}=\mathbf{D}\cdot\mathbf{A}$.
  3. (3)
@@
3.3. Composite Dilated Convolution Layer
Figure 4. Composite Dilated Convolution Framework: After a channel-mixing operation, we split the feature map into multiple parts processed with different dilation rates, and the resulting multi-scale features are concatenated, refined, and upsampled to reconstruct high-resolution segmentation maps.
To expand the receptive field at low cost, we design a Composite Dilated Convolution Layer that processes features in a multi-scale manner. This framework is built upon a cascade of operations that first integrates channel information, then diversifies spatial context via parallel dilated convolutions, and finally refines the combined feature representation.

### v2

@@
Abstract
Semantic segmentation of remote sensing imagery demands precise spatial boundaries and robust intra-class consistency, challenging conventional hierarchical models. To address limitations arising from spatial domain feature fusion and insufficient receptive fields, this paper introduces SAIP-Net, a novel frequency-aware segmentation framework that leverages Spectral Adaptive Information Propagation. SAIP-Net employs adaptive frequency filtering and multi-scale receptive field enhancement to effectively suppress intra-class feature inconsistencies and sharpen boundary lines. Comprehensive experiments demonstrate significant performance improvements over state-of-the-art methods, highlighting the effectiveness of spectral-adaptive strategies combined with expanded receptive fields for remote sensing image segmentation. Our code is available at https://github.com/ZhongtaoWang/SAIP-Net.
keywords:
@@ Abstract
journal: Pattern Recognition
[…]
organization=School of Computer Science, Peking University,addressline=No.5 Yiheyuan Road, Haidian District, city=Beijing, postcode=100871, country=P.R. China
Figure 1: Overview of the challenges and motivations behind SAIP-Net. Remote sensing images often exhibit large intra-class variance, small inter-class differences, and irregular class layouts. To address these challenges, SAIP-Net uses the combination of four modules to improve intra-class consistency and enhance boundary accuracy, significantly improving the segmentation of complex structures in remote sensing images.
1 Introduction
@@ 2 Related Work
Semantic segmentation assigns semantic labels to each pixel and is widely applied in both natural and remote sensing images. While the two domains share foundational techniques, their data characteristics and task constraints differ significantly. Natural image segmentation focuses on perspective-view scenes with diverse objects and consistent contexts [41], while remote sensing imagery features large-scale, top-down views with high intra-class variability, inter-class similarity, which requires domain-specific solutions [42].
2.1 Semantic Segmentation in Natural Images
@@
Early methods, such as thresholding and region growing, were limited by image quality and computational constraints. The advent of deep learning, particularly Convolutional Neural Networks (CNNs) [3] and Fully Convolutional Networks (FCNs) [41], enabled end-to-end pixel-level prediction, forming the basis for encoder-decoder architectures like U-Net [35] and SegNet [36]. These models extract high-level semantic features while preserving spatial resolution.
@@
Subsequent research addressed challenges such as class imbalance, small object detection, and complex backgrounds. Dilated convolutions [8] expanded receptive fields, while PSPNet [9] introduced pyramid pooling to aggregate multi-scale contextual information. Attention mechanisms, such as Coordinate Attention [7], further refine segmentation, especially for small or complex objects. Additionally, unsupervised and semi-supervised approaches [4, 5] have been explored to reduce reliance on large annotated datasets.
2.2 Semantic Segmentation in Remote Sensing Images
@@
Remote sensing segmentation faces unique challenges, including large-scale scenes, spectral diversity, and subtle intra-class variation. While FCN-based methods [11, 12, 13, 14, 15, 16] remain prevalent, additional modules have been developed to enhance spatial and contextual feature extraction. Techniques like channel attention [17] and region-aware methods, such as S-RA-FCN [18] and HMANet [19], address feature redundancy and spatial dependencies. Edge-aware strategies, like Edge-FCN [6], improve boundary precision by incorporating edge detection.
@@
Recently, Transformer-based models [20, 21, 22, 23, 24, 25] have gained attention for their ability to model long-range dependencies via self-attention. For instance, RSSFormer [21] introduces adaptive feature fusion to reduce background noise and enhance object saliency. However, Transformers still face challenges in capturing fine local details and accurately delineating object boundaries.
2.3 Frequency-domain Methods for Semantic Segmentation
@@
Frequency-domain feature processing has emerged as a promising direction in semantic segmentation. Shan et al. decomposed images via the Fourier transform into high and low frequency components [26], which capture edge and body information respectively, ensuring object consistency and edge supervision through deep fusion. Zhang et al. proposed FsaNet [27], leveraging a frequency self-attention mechanism on low-frequency components to drastically reduce computational costs. Chen et al. introduced Frequency-Adaptive Dilated Convolution (FADC) [28], dynamically adjusting dilation rates by local frequency patterns and enhancing feature bandwidth/receptive fields via adaptive kernels and frequency selection. Their FreqFusion [1] further improves boundary sharpness and feature coherence for dense prediction tasks. Ma et al. developed AFANet [29], optimizing semantic structures with a CLIP-guided spatial adapter to boost weakly-supervised few-shot segmentation. These advances underscore the transformative potential of frequency-aware strategies in semantic segmentation.
@@
3 Methods
Figure 2: Overview of the proposed SAIP-Net architecture. The network combines: a Learnable High-Pass Filter Stem to enhance edge details, a Transformer Encoder that extracts global context via multi-stage learnable pooling, and a Frequency-Aware Decoder that fuses high- and low-level features using SAFF modules, while integrated Composite Dilated Convolution Layers expand the receptive field. These modules ultimately lead to result in improved intra-class consistency and enhanced boundary accuracy, thereby significantly improving the segmentation of complex structures in remote sensing images.
@@ 3 Methods
3.1 Transformer Encoder
Our transformer encoder is designed to balance global context modeling and computational efficiency via pooling attention and hierarchical feature extraction. The encoder consists of $4$ stages. Following [2], each stage of our encoder includes multiple Transformer blocks, and spatial downsampling between stages is achieved through learnable pooling operations.
@@ 3.1 Transformer Encoder
$Q=\mathcal{P}_{Q}(X_{in}W_{Q}),\quad K=\mathcal{P}_{K}(X_{in}W_{K}),\quad V=\mathcal{P}_{V}(X_{in}W_{V}),$ (1)
@@
$X_{l}=\mathrm{Attn}(Q,K,V)+Q=\mathrm{Softmax}\left(\frac{QK^{\top}+E^{(\mathrm{rel})}}{\sqrt{D}}\right)V+Q$ (2)
where $E^{(\mathrm{rel})}{ij}=Q_{i}\cdot R{p(i),p(j)}$. Here, $Q_{i}$ represents the query vector at token position $i$, and $R_{p(i),p(j)}$ denotes the learnable relative positional embedding between tokens $i$ and $j$, where $p(i)$ and $p(j)$ map token indices to their spatial locations. The inner product $Q_{i}\cdot R_{p(i),p(j)}$ injects relative position bias into the attention logits. To reduce the number of learnable parameters, we follow [2] and decompose $R_{p(i),p(j)}$ along spatial dimensions as:
$R_{p(i),p(j)}=R^{\mathrm{h}}_{h(i),h(j)}+R^{\mathrm{w}}_{w(i),w(j)},$ (3)
@@
3.2 Spectral Adaptive Feature Fusion Module
Figure 3: (a). Illustration of Spectral Adaptive Feature Fusion (SAFF) module. The module integrates high-level and low-level features using content guided low-pass and high-pass filters alongside spatial offset estimation. The structure of $\mathcal{F}^{\text{HP}}$ is enlarged at the bottom while $\mathcal{F}^{\text{LP}}$ shares a similar architecture. (b). Our module enhances intra-class consistency and boundary accuracy at the feature level, leading to better segmentation results.
@@ 3.2 Spectral Adaptive Feature Fusion Module
Inspired by [1, 30, 31], incorporating a guidance signal (e.g., low-level features) provides valuable local structural priors. These priors help direct the filtering process such that the spatially varying kernels can adapt based on local content, ensuring that essential details like edges and textures are accurately preserved while undesired noise is suppressed. Additionally, guided filters that vary at each position allow fine-grained feature reassembly and precise spatial alignment, ensuring that the fused features accurately reflect both the global context and the local details. Such flexibility directly contributes to improved overall segmentation quality and enhanced boundary delineation.
@@
$\tilde{\mathbf{Y}}^{l+1}=\mathcal{F}^{\text{LP}}(X^{l},\mathbf{Y}^{l+1}),\quad\tilde{\mathbf{X}}^{l}=\mathcal{F}^{\text{HP}}(X^{l},\mathbf{X}^{l})+\mathbf{X}^{l},\quad\mathbf{Z}^{l}=\tilde{\mathbf{Y}}^{l+1}+\tilde{\mathbf{X}}^{l}.$ (4)
@@
$\hat{\mathbf{X}}^{l}=\mathcal{F}^{\text{HP}}(Z^{l},\mathbf{X}^{l})+\mathbf{X}^{l},\quad\hat{\mathbf{Y}}^{l+1}=\mathcal{F}^{\text{LP}}(Z^{l},\mathbf{Y}^{l+1}).$ (5)
@@
$u(i,j),v(i,j)=\mathcal{OG}(Z^{l})_{ij},\quad\mathbf{Y}^{l}_{i,j}=\hat{\mathbf{Y}}^{l+1}_{i+u(i,j),j+v(i,j)}+\hat{\mathbf{X}}^{l}_{i,j}.$ (6)
@@
The key components of this module are two adaptive filter: Content Guided Low-Pass/High-Pass Filter $\mathcal{F}^{\text{LP}}$$\mathcal{F}^{\text{HP}}$ and an offset generator $\mathcal{OG}$. We elaborate on each component as follows:
@@
Inspired by [30, 31], we take $\mathbf{Z}$ as the guidance feature to predict the spatially variant filter weights filtering feature map $\mathbf{X}$. The filter consists of several steps:
  1. 1.
@@
[…]
@@
    $\mathbf{Y}^{g}_{i,j}=\sum_{p,q\in\Omega}\mathbf{\bar{W}}^{g,p,q}_{i,j}\,\mathbf{X}_{i+p,\,j+q},$ (8)
@@
  3. 3.
    Upsampling of High-Level Features. If the input is high-level features, the four groups $\mathbf{Y^{g}}$ are rearranged using Pixel Shuffle[32] to obtain a smooth $2\times$ upsampled feature map $\mathbf{Y}\in\mathbb{R}^{C\times 2H\times 2W}$.
@@
    $\mathbf{S}^{p,q}_{i,j}=\frac{\sum_{c}\mathbf{Z}_{c,i,j}\,\mathbf{Z}_{c,i+p,j+q}}{\sqrt{\sum_{c}(\mathbf{Z}_{c,i,j})^{2}}\;\sqrt{\sum_{c}(\mathbf{Z}_{c,i+p,j+q})^{2}}},$ (9)
  2. 2.
@@
    $\mathbf{D}=\text{Conv}_{3\times 3}([\mathbf{Z},\mathbf{S}]),\quad\mathbf{A}=\sigma\!\bigl(\text{Conv}_{3\times 3}([\mathbf{Z},\mathbf{S}])\bigr),$ (10)
    Where [][...] denotes the concat operation and $\sigma$ denotes the sigmoid function. Finally, the offset is given by $\mathbf{O}=\mathbf{D}\cdot\mathbf{A}$.
  3. 3.
@@
3.3 Composite Dilated Convolution Layer
Figure 4: Composite Dilated Convolution Framework: After a channel-mixing operation, we split the feature map into multiple parts processed with different dilation rates, and the resulting multi-scale features are concatenated, refined, and upsampled to reconstruct high-resolution segmentation maps.
To expand the receptive field at low cost, we design a Composite Dilated Convolution Layer that processes features in a multi-scale manner. As shown in Figure 4, this framework is built upon a cascade of operations that first integrates channel information, then diversifies spatial context via parallel dilated convolutions, and finally refines the combined feature representation.